OpenStudy (anonymous):

How do you put the points (-3,0) and (0,6) into a standard form?

OpenStudy (dan815):

what is slope of the line that these 2 points are on

OpenStudy (anonymous):

use the slope formula to figure out the slope. then after that you can use the point-slope formula to get to standard form :D (ax+by=c)

OpenStudy (texaschic101):

do you know how to find the slope ?

OpenStudy (anonymous):

ok, now I need help solving: y-2=1/2x(x+5)

OpenStudy (anonymous):

@eohmeo @texaschic101

OpenStudy (texaschic101):

did you get the first problem ? Now you need help with the other one ?

OpenStudy (texaschic101):

y - 2 = 1/2(x + 5) Is this your problem ?

OpenStudy (anonymous):

no I didn't figure out the other problem (I forgot to add "the x-intercept is 6 and the y-intercept is 3" and then I changed them into points" in the question. maybe you can also help me with this one? @texaschic101

OpenStudy (anonymous):

yes y-2=1/2(x+5) is the one im working on

OpenStudy (texaschic101):

y - 2 = 1/2(x + 5) y - 2 = 1/2x + 5/2 y = 1/2x + 5/2 + 2 y = 1/2x + 5/2 + 4/2 y = 1/2x + 9/2

OpenStudy (texaschic101):

you need me to do the first problem ?

OpenStudy (anonymous):

yes please! thanks so much

OpenStudy (anonymous):

oh wait for the y=1/2x+9/2, how did you add two to both sides and add 4/2 to 5/2?

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-3,0) x1 = -3 and y1 = 0 (0,6) x2 = 0 and y2 = 6 now lets sub slope(m) = (6 - 0) / (0 -(-3) m = 6/3 m = 2 now use y - y1 = m(x - x1) slope(m) = 2 (-3,0) x1 = -3 and y1 = 0 now sub y - (-3) = 2(x - 0) y + 3 = 2x - 0 y = 2x - 3 standard form is Ax + By = C y = 2x - 3 -2x + y = -3 (multiply by -1 to make x positive) 2x - y = 3 <== standard form

OpenStudy (texaschic101):

y - 2 = 1/2(x + 5) -- distribute through the parenthesis y - 2 = 1/2x + 5/2 -- now we will add 2 to both sides to get y by itself y - 2 + 2 = 1/2x + 5/2 + 2 -- combine like terms y = 1/2x + 5/2 + 2 -- need common denominator which is 2 y = 1/2x + 5/2 + 4/2 y = 1/2x + 9/2 understand ?

OpenStudy (texaschic101):

do you need that one in standard form also ?

OpenStudy (anonymous):

oh yes! thank you:) i understand that one, but i don't understand a part of the (-3,0) and (0,6) one... 0 is y1 not -3, so shouldn't it be y-0=2(x+3)?

OpenStudy (anonymous):

no i don't need it in standard form, but thanks

OpenStudy (texaschic101):

crap...your right...let me do it again y - 0 = 2(x + 3) y - 0 = 2x + 6 y = 2x + 6 Ax + By = C -2x + y = 6 (multiply by -1 to make x positive) 2x - y = -6 Thanks for catching my mistake :)

OpenStudy (anonymous):

heheh thank you! ;D

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

yes you helped A LOT! I understand everything:D K-bye!